OpenStudy (anonymous):

What is the point-slope form of the equation of a line that passes through the point (6, −8) and has a slope of −2?

OpenStudy (anonymous):

im confused about this problem can u help me more

pooja195 (pooja195):

\[y−(-8)=-2(x−6)\]

pooja195 (pooja195):

refresh your page

pooja195 (pooja195):

now solve the problem from here

OpenStudy (anonymous):

i did

pooja195 (pooja195):

ok now solve the problem

OpenStudy (anonymous):

the thing is i dont even know how to solve it i keep on trying but it just dont make sense to me

pooja195 (pooja195):

\[y−(−8)=−2(x−6)\] 1) -(-8)=+8 so it would be y+8 2) distribute -2(x-6) 3) -2x+12 4) y+8=-2x+12 <----solve from here

OpenStudy (anonymous):

dont i have to subtract 8 from both sides or somthing

OpenStudy (anonymous):

are u there

pooja195 (pooja195):

yes i asked you to solve it from there

OpenStudy (anonymous):

is the answer y-(-8)=-2x+6

OpenStudy (anonymous):

can someone please tell me the answe

pooja195 (pooja195):

You want the answer this is not the place for answers we dont give answers thats not how it works ive literally done all the solving for you you just got to do a few more steps

pooja195 (pooja195):

4) y+8=-2x+12 <----solve from here

pooja195 (pooja195):

whats 12-8?

OpenStudy (anonymous):

4

pooja195 (pooja195):

y=2x+4 <----answer

OpenStudy (anonymous):

solve that and i get my answer

pooja195 (pooja195):

y=2x+4 <----ANSWER hint: that is your answer we r done with the problem the end

OpenStudy (anonymous):

but thats not on my multiple choice question

OpenStudy (anonymous):

help please